OpenStudy (anonymous):

dy/dx of 6x^2+3xy+2y^2=-17y+6

OpenStudy (anonymous):

\[6x^2+3xy+2y^2=-17y+6\] diff. w.r.t. x \[12x+3\left( x \frac{ dy }{ dx }+y*1 \right)+4y \frac{ dy }{ dx }=-17\frac{ dy }{ dx }\] simplify and find dy/dx

OpenStudy (haseeb96):